OpenStudy (anonymous):

(4 - square root of 3) (12 + 5 square root of 3)

OpenStudy (anonymous):

Verify. \[(4 - \sqrt{3}) (12 + 5 \sqrt{3})\]

OpenStudy (anonymous):

\[33+8\sqrt{3}\]

OpenStudy (angela210793):